• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name tajima

    Log entry time 16:12:35 on January27,2002

    Entry number 81026

    keyword=target ioc reboot-- PID problem

    At15:30, I noticed that the loop3 temperature is bit too high(19.80 from TV camera). I talked JP and he thinks the PID was not working and told me to reboot IOC.
    I did it and now everything looks ok.